OpenStudy (anonymous):

does anyone on know how to solve by elimination I posted -9-9x-4y=0 -11x-11-11y=0

OpenStudy (anonymous):

Set them = to each other

OpenStudy (anonymous):

-9x-4y-9=0 -11x-11-11y=0?

OpenStudy (asnaseer):

divide the 2nd equation by 11 to get: -x -1 -y = 0 then get an expression for x from this by adding x to both side: -x -1 -y + x = 0 + x -1 -y = x now substitute this expression for x into the first equation and solve.

OpenStudy (campbell_st):

rewite the equations as 11y = -11x -11 then y = (-x -1) sub into the 1st equation -9x -4(-x -1) = 9 then solve for x so I think x = -1 and y = 0

OpenStudy (anonymous):

I don't understand

OpenStudy (asnaseer):

you are given two equations with two unknowns in each equation (x and y). the idea is to eliminate one of the unknowns and then solving the resulting equation with just one unknown.

OpenStudy (anonymous):

i was told you either add or subtract from one side

OpenStudy (asnaseer):

you can also multiply or divide both sides by some number

OpenStudy (asnaseer):

your 2nd equation had a common factor of 11 in every term - so if you divide by 11, you get a much simpler equation